Text Size

02-13-2010
February 13, 2010
ISS On-Orbit Status 02/13/10

All ISS systems continue to function nominally, except those noted previously or below. Saturday – a busy “day” for the combined crews. Sleep shift in effect: The ISS crew’s workday began last evening at 4:14pm and ended this morning at 7:44 EST (see time table at bottom).

At wake-up last evening, FE-1 Suraev did the regular daily early-morning check of the aerosol filters at the Russian Elektron O2 generator which Maxim had installed on 10/19 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). [FE-1 again inspected the filters this morning before bedtime, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]

CDR Williams & FE-6 Creamer continued their new week-long session of the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight), TJ’s second, donning their Actiwatches, from which to log data to the HRF-1 (Human Research Facility 1) laptop. [To monitor the crewmembers’ sleep/wake patterns and light exposure, the crewmembers sometimes wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

Working in the SM (Service Module), after switching the Vozdukh CO2 removal system to Automatic mode, Kotov & Suraev finished the installation of a new BRI Smart Switch Router computer and ASP Network Connection Adapter, started 2/11, laying new cables from the BRI to the old BRI location behind the medical locker and from the new BRI to the RSS1 and RSS2 laptops, plus from the ASP to future payload Ethernet connections behind panel 225. Later, Oleg switched Vozdukh back to Manual mode.

Maxim again downloaded TEKh-15/DAKON-M IZGIB (“Bend”) structural dynamics measurements of the 20A docking and re-set the hardware for more data taking, later checking its status. [IZGIB has the objective to help update mathematical models of the ISS gravitation environment, using accelerometers of the Russian SBI Onboard Measurement System, the GIVUS high-accuracy angular rate vector gyrometer of the SUDN Motion Control & Navigation System and other accelerometers for unattended measurement of micro-accelerations at science hardware accommodation locations - (1) in operation of onboard equipment having rotating parts (gyrodynes, fans), (2) when establishing and keeping various ISS attitude modes, and (3) when performing crew egresses into space and physical exercises.]

CDR Williams replaced a failed LHA (Lamp Housing Assembly) in the starboard CQ (Crew Quarters), stowing the old unit in a CTB (Cargo Transfer Bag) in COL (Columbus Orbital Laboratory).

Afterwards, Jeff filled the WRS WSTA (Water Recovery System / Waste Storage Tank Assembly) to start UPA (Urine Processor Assembly) processing. The filling was later repeated by FE-6 Creamer.
[During the first processing run, the WPA MLS (Water Processor Assembly / Mostly Liquid Separator) inlet pressure dropped from ~50 mmHg (gauge) to 8 mmHg (gauge), indicative of a clogged EFA filter or solenoid valve. Processing completed despite the low pressures, as did a second processing run. The current plan is to continue processing in the degraded mode, with the priority given to processing of urine distillate. It is expected that starting the WPA processing cycle at a higher tank quantity will allow the MLS to fill and cycle, which causes a "reverse flush" of the WWT (Waste Water Tank) solenoid valve. This should provide sufficient data to isolate the observed pressure drop to either the solenoid valve or the recently-installed EFA filter assembly. If the problem is confirmed to be the EFA, the team will pursue the plan to flush the WWT and replace the EFA.]

After completing the fine leak check of Node-3 “Tranquility” last night, Williams & MS2 Steve Robinson opened the Node-3 hatch (~9:18pm EST) and began the lengthy job of outfitting the Node-3 vestibule, first moving equipment from Node-1 to the vestibule to gain ingress into “Tranquility” (docked to Node-1), then taking air sample readings in Node-3 with the GRC (Grab Sample Container) and CSA-CP (Compound Specific Analyzer-Combustion Products).

Jeff & Steve then mounted caps on the Node-3 aft & nadir PPRVs (Positive Pressure Relief Valves), and installed data, fluid and IMV (Intermodular Ventilation) jumpers in the Node-1-to-Node-3 vestibule.

Next step was for Jeff to configure Node-3 rack bays to accept their racks, followed by the transfer of the ARS (Atmosphere Revitalization Systems) Rack from Kibo JPM (JEM Pressurized Module, loc. F3) by Soichi Noguchi, assisted by PLT Terry Virts, to Node-3 (loc. A4). Soichi also removed the temporary IMV duct from Node-2 and installed it in Node-3. [These tasks took longer than expected, and Jeff was allowed to waive his exercise sessions today.]

TJ Creamer took airflow measurements with the Velocicalc instrument at the Node-3’s temporary air duct to verify sufficient airflow, then conducted an inspection of the aft NPRVs (Negative Pressure Relief Valves) for actuation during launch. [Node-3 Stbd NPRVs will be inspected during IMV valve Installation.]

Afterwards, TJ conducted purges and leak checks on the Node-3 O2 and N2 (nitrogen) systems.

Creamer then installed HYZOD window covers and protective shields in Node-3. [This included the outer cover assembly on the Stbd window plus inner covers on the Node-3 Zenith & Stbd windows, also covers on the EVA hatch and IVA hatch windows, as well as protective shields on Node-3 Nadir, Fwd, Aft & Port.]

After collecting plant samples (wheat) of the Russian BIO-5 Rasteniya-2 ("Plants-2") experiment for return to the ground, Max Suraev transferred new Rasteniya-2 hardware from Progress 36P and installed it in the SM, setting up the root module, plants kit, leaf chamber, light unit and KDV water container, then filling two water tank, loading new software and running a hardware test, then starting the experiment. [Rasteniya-2 researches growth and development of plants under spaceflight conditions in the Lada-15 greenhouse from IBMP (Institute of Bio-Medical Problems, Russian: IMBP). During its operation, the experiment requires regular daily maintenance of the experiment involving monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, and photo/video recording.]

Working with FE-4 Kotov, Noguchi transferred the ARED Advanced Resistive Exercise Device) to the Node-3.

After stowing the ARED advanced resistive exercise device overnight in Node-3 and re-installing the
VIS (Vibration Isolation Stabilization) in Node-3, Suraev & Noguchi today installed the ARED in “Tranquility”, including all cables, and taking documentary photography. [Upon completion of ARED installation, the ground will review the photos and provide a Go/No Go for the unmanned ACO (Activation & Checkout), scheduled with the Manned ACOs on FD9 (2/16).]

Maxim started charging the Sony SM-HDV-U03 battery for tomorrow’s session of the Russian payload TkhN-7 SVS (Self-Propagating High-Temperature Synthesis). [SVS uses its own camera, “Telescience” hardware from PK-3 (Plasma Crystallization) and the onboard Klest TV system for researching self-propagating high-temperature fusion of samples in space.]

Terry & Kay operated the SSRMS (Space Station Remote Manipulator System) to release Node-3 and grapple the Lab, followed by a base change from Node-2 to the Lab. Virts & Hire then released Node-2 and maneuvered the Canadian arm to the overnight park position. [The SSRMS remained there overnight and then supported the Cupola and Thermal Radiator Rotary Joint (TRRJ) photogrammetry activity scheduled to begin later this morning.]

The crewmembers worked out with their regular 2-hr physical exercise on the CEVIS cycle ergometer (FE-5, FE-6), TVIS treadmill (FE-1, FE-4), T2/COLBERT advanced treadmill (FE-1, FE-5, FE-6), and VELO bike ergometer with bungee cord load trainer (FE-4).

At ~4:09am-5:09am EST this morning, the joint crew held an in-depth one-hour review of procedures for the EVA-2 spacewalk, with egress scheduled tonight at ~9:14pm EST.

At ~6:39am, Behnken (EV1) & Patrick (EV2) began their “campout” (nachalo desaturatsiy = desaturation start) in the A/L with hatch closure and depressurization of the CL (Crewlock) from 14.7 to 10.2 psi, followed by mask prebreathe (~6:39am-7:44am) and sleep from 8:14am-4:14pm tonight. A hygiene break is scheduled for 4:54pm-6:04pm. EVA-2 objectives: Node-3 outfitting (hooking up ammonia lines), support grappling Cupola. [Sleep for the ISS crew began at 7:44am.]

Cupola Issue: The crew could not install the Node-3 Axial Port Center Disk Cover due to physical interference with Cupola structure such as a Cupola standoff bar. Teams are investigating center disk options such as removal of Cupola structure, alternate installation of cover, etc., and Cupola relocation options such as EVA removal of Node-3 nadir cover, delay of Cupola relocation, etc. Cupola grapple and relocate are expected to happen on schedule.

Orbiter TPS: The Shuttle Thermal Protection System has been cleared for entry.

  • Sleep shifting started with the Progress docking on 2/4. On 2/6, crew wake shifted earlier, to 5:40pm EST. 20A Undock will drive Crew Wake one and a half hours earlier to 4:09pm by FD12. This shift is accomplished by moving Crew Sleep 30 min earlier on FDs 6-10, and then again on FD11 and FD12. Wake/Sleep table:

FD
Cal
GMT
WAKE
EST
SLEEP
EST
6
13-Feb
44
4:14pm (2/12)
7:44am
7
14-Feb
45
4:14pm (2/13)
7:44am
8
15-Feb
46
4:14pm (2/14)
7:44am
9
16-Feb
47
4:14pm (2/15)
7:44am
10
17-Feb
48
4:14pm (2/16)
7:44am
11
18-Feb
49
4:14pm (2/17)
6:44am
12
19-Feb
50
3:14pm (2/18)
6:44am
13
20-Feb
51
1:00am
4:30pm


Weekly Science Update (Expedition Twenty-Two -- Week 11)

3-D SPACE: No report.

AgCam (Agricultural Camera): No report.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ALTEA DOSI (NASA/ASI): No report.

APEX (Advanced Plant Experiments on Orbit) -Cambium: No report.

APEX-TAGES (Transgenic Arabidopsis Gene Expression System): No report.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): No report.

BIOLAB (ESA): No report.

BIORHYTHMS (JAXA, Biological Rhythms): No report.

BISE (CSA, Bodies in the Space Environment): No report.

BISPHOSPHONATES: “Jeff & Soichi, thanks for completing your pill ingestion. Your next session is scheduled for 2/14.”

CARD (Long Term Microgravity: Model for Investigating Mechanisms of Heart Disease, ESA): No report.

CARDIOCOG-2: Complete.

CB (JAXA Clean Bench): No report.

CBEF (JAXA Cell Biology Experiment Facility)/SPACE SEED: CBEF fan repair (micro-G section) on 1/29.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): No report.

CERISE (JAXA): Samples will return on the next shuttle.

CETSOL (ESA): 2 processed sample cartridge assemblies (SCAs) will be downloaded on 20A.

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Ongoing.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack), MDCA/Flex: No report.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial Photo (JAXA): No report.
CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DECLIC (Device for the Study of Critical Liquids & Crystallization, CNES/NASA): No report.

DomeGene (JAXA): Complete.

DOSIS (ESA): No report.

EarthKAM (Earth Knowledge Acquired by Middle School Students): No report.

EDR (European Drawer Rack, ESA): No report.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EMCS (European Modular Cultivation System): No report.

ENose (Electronic Nose): No report.

EPM (European Physiology Module): No report.

EPO (Educational Payload Operations, NASA): No report.

EPO LES-2 (ESA): No report.

EPO 3-min Video (JAXA): No report.

EPO J-Astro Report (JAXA): Ongoing.

EPO Dewey’s Forest (JAXA): Watering in progress.

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): No report.

EPO Moon Score (JAXA): No report.

EPO Try Zero-G (JAXA): No report.

EPO Kibo Kids Tour (JAXA): Complete.

EPO Paper Craft (Origami, JAXA): Soichi made origami on 2/8.

EPO Poem (JAXA): No report.

EPO Spiral Top (JAXA): No report.

ETD (Eye Tracking Device): Completed.

FACET (JAXA): No report.

FIR/LMM/CVB (Fluids Integrated Rack / Light Microscopy Module / Constrained Vapor Bubble): No report.

FWED (Flywheel Exercise Device, ESA): No report.

FOAM STABILITY (ESA): Tapes were returned on ULF3.

FOCUS: “Thank you Jeff for performing the FOCUS experiment. The images have been received on the ground.”

FSL (Fluid Science Laboratory, ESA): No report.

GEOFLOW: No report.

HAIR (JAXA): Soichi collected hair sample from TJ on 1/19. “Thank you for bearing the pain.”

HDTV System (JAXA): Was delivered by HTV1.

Holter ECG (JAXA): No report.

HQPC (JAXA): Was delivered by 34P.

HREP (HICO/Hyperspectral Imager for the Coastal Ocean & RAIDS/Remote Atmospheric & Ionospheric Detection System/JAXA): Due to power constraints, HREP will be in low power mode from 2/15 until 2/16..

ICE CRYSTAL (JAXA): Complete.

ICV (Integrated Cardiovascular): No report.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

INTEGRATED IMMUNE: No report.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): No report.

IRIS (Image Reversal in Space, CSA): No report.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): No report.

Marangoni UVP (JAXA): No report.

MAXI (Monitor of All-sky X-ray Image, JAXA): Continuing observation operation.

Marangoni UVP: Going well. Oscillation flow is under observation.

MEIS (Marangoni Experiment for ISS) in JAXA FPEF (Fluid Physics Experiment Facility): No report.

MDCA/Flex: See under CIR.

Microbe-1 (JAXA): No report.

Micro-G Clay (JAXA EPO): Complete.

MMA (JAXA/Microgravity Measurement Apparatus): No report.

MISSE7 (Materials ISS Experiment): To prevent overheating during the ISS rotation and shuttle docking, MISSE7 powered down PEC 7b and 7a nadir on Feb 9. Temperatures will be monitored during the docked period and if thermal conditions are acceptable the PECs will be returned to normal operations.

MSG-SAME (Microgravity Science Glovebox): Complete.

MSL (Materials Science Laboratory): MSRR-1 activities have resumed on 1/20, with the processing of the CETSOL#3 Sample Cartridge Assembly. MICAST: Solidification phase is planned to take place during 1/22 night time. 2 processed sample cartridge assemblies (SCAs) will be downloaded on 20A. Next SCA to be processed is MICAST#2.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC-1 “Pirs”.

MULTIGEN-1: Completed.

NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): No report.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY/ProK: No report.

PADLES (JAXA, Area PADLES 3; Passive Area Dosimeter for Lifescience Experiment in Space): No report.

PASSAGES (JAXA): No report.

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): No report.

PCG (JAXA, Protein Crystal Growth): No report.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

POLCA/GRAVIGEN (ESA): Complete.

RadGene & LOH (JAXA): Complete.

RadSilk (JAXA): Samples were returned to ground on ULF3.

RST/Reaction Self Test (Psychomotor Vigilance Self Test on the ISS): “Jeff, we have provided a pdf with the feedback you requested regarding your data. We expect you should receive it soon. Please let us know if you have any questions. Jeff, TJ, and Soichi: thank you for your participation in Reaction Self Test!”

SAIBO Rack (JAXA): “CB port cap reattached, 1/28.”

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): No report.

SAMPLE: Complete.

SCOF (Solution Crystallization Observation Facility, JAXA): No report.

SEDA-AP (Space Environment Data Acquisition Equipment-Attached Payload, JAXA): Continuing observation operations.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLAMMD (Space Linear Acceleration Mass Measurement Device): “Jeff and TJ, thank you for your continued support of SLAMMD. As mentioned in the daily summary before your last SLAMMD BMM, a new software version for the hardware was uplinked and we appreciate you calling down with body mass data observations. After analyzing the data, we have requested that a SLAMMD calibration activity be scheduled to support this software update. Until this can be accommodated in the timeline, the ground team will continue to recalculate the measurements with the appropriate algorithm. The team looks forward to receiving each session’s BMM data, as it is invaluable feedback as they tweak the SLAMMD software performance.”

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Jeff: Thanks for completing the Actiwatch download and changing the 9V battery in the Actiwatch Reader. Jeff and TJ: You are currently completing another week of Sleep logging.”

SMILES (JAXA): Continuing observations.

SODI/IVIDIL (Selectable Optical Diagnostics Instrument/Influence of Vibration on Diffusion in. Liquids, ESA): The Cell Array #2 has a binary mixture of 50% water - 50% isopropanol leading to positive Soret coefficient (when the lighter component is driven towards the higher temperature region of the experimental cell). IVIDIL remaining runs have been successfully performed from 1/11 until 1/20 on a 24/7 basis. This last script concluded a successful set of 54 runs, and the experiment is now completed. Preliminary assessment of science images on-going.

SOLAR (Solar Monitoring Observatory) Sun observation window #25 started on 2/9. The tracking windows on each orbit have been shorter than usual due to the attitude of the ISS.

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): No report.

SPICE (Smoke Point In Co-flow Experiment): No report.:

SPINAL (Spinal Elongation): | No report.

SWAB (Characterization of Microorganisms & Allergens in Spacecraft): “Jeff, thank you for completing the 7th of 8 SWAB Water Collections! Only one more to go!”

THERMOLAB (ESA): “Thanks Jeff for your fourth THERMOLAB session. The data has been provided to the science team.”

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

VASCULAR (CSA): “Jeff, thanks for completing your blood draw. The PI team will be in Russia to welcome you back and complete post-flight BDC with you.”


VO2max (NASA): No report.

VLE (Video Lessons ESA): VLE-1 completed.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): No report.

YEAST B (ESA): No report.

CEO (Crew Earth Observations): No report.

No CEO (Crew Earth Observation) photo targets uplinked for today.

Significant Events Ahead (all dates Eastern Time and subject to change):
02/10/10 -- STS-130/Endeavour/20A docking (12:57am)
· 02/13/10 -- EVA-2 (9:14pm; Node-3 outfitting), Node-3 activation
· 02/14/10 -- Cupola relocation & outfitting
· 02/16/10 -- EVA-3 (9:14pm; Node-3, Cupola, and PMA3 tasks)
02/17/10 -- Final cargo transfers, hatch closure, reboost
02/18/10 -- STS-130/Endeavour/20A undock (7:35pm)
02/20/10 -- STS-130/Endeavour/20A KSC deorbit burn (8:59pm)
02/20/10 -- STS-130/Endeavour/20A KSC landing (10:01pm)
03/18/10 -- Soyuz TMA-16/20S undock/landing (4:10am/7:30am EST)
--------------Three-crew operations-------------
03/18/10 -- STS-131/Discovery/19A – MPLM(P), LMC (launch ~1:30pm EST)
04/02/10 -- Soyuz TMA-18/22S launch – Skvortsov (CDR-24)/Caldwell/Kornienko
04/04/10 -- Soyuz TMA-18/22S docking
--------------Six-crew operations-----------------
04/27/10 -- Progress M-03M/35P undock
04/28/10 -- Progress M-05M/37P launch
04/30/10 -- Progress M-05M/37P docking
05/14/10 -- STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1 (~2:00pm EST)
05/10/10 -- Progress M-04M/36P undock
05/31/10 -- Soyuz TMA-17/21S undock/landing
--------------Three-crew operations-------------
06/14/10 -- Soyuz TMA-19/23S launch – Wheelock (CDR-25)/Walker/Yurchikhin
06/16/10 -- Soyuz TMA-19/23S docking
--------------Six-crew operations-----------------
07/xx/10 -- US EVA-15
07/xx/10 -- Russian EVA-25
06/28/10 -- Progress M-06M/38P launch
07/02/10 -- Progress M-06M/38P docking
07/26/10 -- Progress M-05M/37P undock
07/27/10 -- Progress M-07M/39P launch
07/29/10 -- Progress M-07M/39P docking
07/29/10 -- STS-134/Endeavour (ULF6 – ELC3, AMS-02) (~7:30am EST)
08/30/10 -- Progress M-06M/38P undock
08/31/10 -- Progress M-08M/40P launch
09/02/10 -- Progress M-08M/40P docking
09/15/10 -- Soyuz TMA-18/22S undock/landing
09/16/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) (~12:01pm EST)
09/18/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) docking
09/22/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) undock
09/30/10 -- Soyuz TMA-20/24S launch – Kelly (CDR-26)/Kaleri/Skripochka
10/xx/10 -- Russian EVA-26
10/26/10 -- Progress M-07M/39P undock
10/27/10 -- Progress M-09M/41P launch
10/29/10 -- Progress M-09M/41P docking
11/15/10 -- Soyuz TMA-19/23S undock/landing
11/30/10 -- Soyuz TMA-21/25S launch – Kondratyev (CDR-27)/Coleman/Nespoli
12/15/10 -- Progress M-08M/40P undock
02/08/11 -- Progress M-09M/41P undock
02/09/11 -- Progress M-10M/42P launch
02/11/11 -- Progress M-10M/42P docking
03/30/11 -- Soyuz TMA-22/26S launch
xx/xx/11 -- Progress M-11M/43P launch
05/30/11 -- Soyuz TMA-23/27S launch
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton.